OpenStudy (anonymous):

well heres another chock full of questions: what is gfc for: 60,45,24,72,32,48,100,30,54,36,3,97,5,4,20,24,36,9,45 and yes im horrible at this.

OpenStudy (lilg132):

put it in order first

OpenStudy (matt101):

There isn't a greatest common factor for all of them...e.g. 4 and 45 have no factors in common... Are they supposed to be in pairs?

OpenStudy (anonymous):

yes

OpenStudy (matt101):

45 = 3 * 3 * 5 60 = 2 * 2 * 3 * 5 ------------------ GCF= 3 * 5 = 15 LCM=2 * 2 * 3 * 3 * 5 = 180 Try the others like that :)

OpenStudy (anonymous):

(60,45) (24,72) (32,48) (100,30) (54,36) (3,97,5) (4,20,24) (36,9,45) those are the pairs.

OpenStudy (jamesj):

matt101 has given you a good algorithm. Try it on the pair (24,72) and see what you get.

OpenStudy (anonymous):

OMG I totally nessed u up last time. I gave u the lcm and noone corrected me

OpenStudy (anonymous):

Hey I am sorry so totally ignore the answer I gave u last time

OpenStudy (anonymous):

I will go to the previous post and correct it sorry

OpenStudy (anonymous):

ok

OpenStudy (matt101):

To find GCF or LCM, first break down your numbers into prime factors, and organize it as I did above. If both numbers share the same factor, put it in the same column, otherwise give the factor its own column. Multiply all the factors that are shared between the two numbers to get the GCF, multiply the numbers from every column to get the LCM. Try the method above for 24 and 72 and see if you get it.

OpenStudy (jamesj):

The prime factorization of 24 is 24 = 2.2.2.3, where . means multiplication. You find the factorization now for 72 and compare it to 24.

OpenStudy (anonymous):

im trying the multiples of 4 to get to 72

OpenStudy (jamesj):

It might be easier to break it down as 72 = 8 x 9 Now 8 =2x2x2 and 9 = 3x3 Hence 72 = ... what?

OpenStudy (anonymous):

im in 7th just so you know

OpenStudy (anonymous):

I gave you the correct answer for that previous one I did give u the gcf

OpenStudy (anonymous):

Won't the gcf for all these numbers be 1?

OpenStudy (jamesj):

well, you do know that 72 = 8 x 9 yes? Well if 72 = 8x9 and 8 = 2x2x2 and 9=3x3, it must be that 72 = 8 x 9 = 2x2x2 x 3x3

OpenStudy (anonymous):

Is that correct GCF?

OpenStudy (jamesj):

@rld613 we're not finding the gcf of all the numbers. But yes, it would be if that were the question.

OpenStudy (anonymous):

So what is the question asking?

OpenStudy (jamesj):

Now as 24 = 2x2x2x3 72 = 2x2x2x3x3 what is the gcf(24,72)? Look at what matt101 did above.

OpenStudy (jamesj):

@rld613: read the thread. It turns out the problem is the find the gcf of a number of pairs and triples of integers. This example, gcf(24,72) is the second problem.

OpenStudy (anonymous):

Oh I see

OpenStudy (anonymous):

U cld do it the long way like I showed u b4hand

OpenStudy (anonymous):

like take every pair and find the factors of each one

OpenStudy (anonymous):

i dont understand jamesj. what are you getting at here?

OpenStudy (jamesj):

Here is what matt101 did with the first example: 45 = 3 * 3 * 5 60 = 2 * 2 * 3 * 5 ------------------ GCF= 3 * 5 = 15

OpenStudy (anonymous):

Do you know how to do the factor tree?

OpenStudy (matt101):

Check out what I did above. fourth post down

OpenStudy (jamesj):

He found the prime factors in common across both 45 and 60. The product of those factors is the gcf.

OpenStudy (matt101):

i.e. what James just reposted haha

OpenStudy (anonymous):

Solike how did ur teacher teach u to find the gcf?

OpenStudy (jamesj):

So now, here again is the example for 60 and 45 45 = 3 * 3 * 5 60 = 2 * 2 * 3 * 5 ------------------ GCF= 3 * 5 = 15 Apply that to 24 and 72: 24 = 2 x 2 x 2 x 3 72 = 2 x 2 x 2 x 3 x 3 ------------------- GCF =

OpenStudy (anonymous):

she didnt

OpenStudy (anonymous):

ok so basically when I taught it 8th graders i taught it like this

OpenStudy (anonymous):

So lets say i took the number 45

OpenStudy (jamesj):

@rld613: you're not helping here. We have an algorithm. Unless it is wrong, please stop posting for a minute and let him work the problem.

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

But he is in 7th grade that is very young

OpenStudy (anonymous):

Y don'y u teach him the long way

OpenStudy (jamesj):

Let him try for heaven's sake. For the last time, here again is an example, and here is our problem: So now, here again is the example for 60 and 45 45 = 3 * 3 * 5 60 = 2 * 2 * 3 * 5 ------------------ GCF= 3 * 5 = 15 Apply that to 24 and 72: 24 = 2 x 2 x 2 x 3 72 = 2 x 2 x 2 x 3 x 3 ------------------- GCF = What are the factors in common between 24 and 72?

OpenStudy (anonymous):

oh so then isnt it that the gcf is 2 and 3?

OpenStudy (anonymous):

there can only be one answer what is the greatest common factor remembeer greatest

OpenStudy (anonymous):

oh its 3

OpenStudy (anonymous):

Neways sorry for bothering I am gonna go

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

yup that is the answer

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

James doesn't think so

OpenStudy (matt101):

No it isn't...

OpenStudy (jamesj):

24 = 2 x 2 x 2 x 3 72 = 2 x 2 x 2 x 3 x 3 ------------------- GCF = 2 x 2 x 2 x 3 The factors in common are 2 and 2 and 2, and 3. So the gcf(24,72) = 2x2x2x3 = 8x3 = 24

OpenStudy (anonymous):

Oh i see what you are doing. I didn't get it

OpenStudy (anonymous):

Oh i like this method

OpenStudy (jamesj):

Just like the factors in common between 45 and 60 are 3 and 5: 45 = 3 * 3 * 5 60 = 2 * 2 * 3 * 5 ------------------ GCF= 3 * 5 = 15

OpenStudy (jamesj):

Hence gcf(45,60) = 3x5 = 15.

OpenStudy (anonymous):

so then for 60 and 45 it would be 15

OpenStudy (jamesj):

Yes. Try it on your next example: (32,48) =========== Remaining problems: (100,30) (54,36) (3,97,5) (4,20,24) (36,9,45)

OpenStudy (jamesj):

First write out the prime factorization 32 = what? 48 = what?

OpenStudy (anonymous):

yeah my computer does that alot

OpenStudy (anonymous):

32= 4 x 8

OpenStudy (jamesj):

Write that as the product of prime numbers. 4 and 8 aren't prime.

OpenStudy (anonymous):

4= 2 x 2 8= 4 x 4 right?

OpenStudy (jamesj):

no

OpenStudy (anonymous):

so then 2 x 2 x 2= 8

OpenStudy (jamesj):

yes, hence 32 = ...

OpenStudy (anonymous):

i meant to put another 2

OpenStudy (jamesj):

What you wrote last was correct. 8 = 2 x 2 x 2, and 4 = 2 x 2. Hence 32 = 8 x 4 = 2x2x2 x 2x2 Now you factor 48

OpenStudy (jamesj):

hint: 48 = 6 x 8

OpenStudy (anonymous):

so then i have to do the factors of 2 so that i can get to 6

OpenStudy (jamesj):

No. You find whatever the prime factorization of 6 is, whatever it is.

OpenStudy (jamesj):

6 = 2 x ...what?

OpenStudy (anonymous):

which is 2x3 right?

OpenStudy (jamesj):

of course. Hence the prime factorization of 48 is ... what? 48 = ...

OpenStudy (jamesj):

48 = 8 x 6 = ...

OpenStudy (anonymous):

i think i messed up

OpenStudy (jamesj):

8 = 2x2x2 and 6 = 2x3 Hence 48 = 8 x 6 = ...what?

OpenStudy (anonymous):

2

OpenStudy (jamesj):

what do you mean. 48 does not equal 2. Try that again.

OpenStudy (jamesj):

8 = 2x2x2 and 6 = 2x3 Hence 48 = 8 x 6 = ...what?

OpenStudy (matt101):

No you're on the right track. You got the prime factorization of 6, which is 2x3. You found the prime factorization of 8 already, which was 2x2x2. Therefore, the prime factorization of 48, which is 8x6, is what?

OpenStudy (anonymous):

no i mean what the common factor is

OpenStudy (anonymous):

2

OpenStudy (jamesj):

No, that's wrong. Keep with the method here. What is the prime factorization of 48

OpenStudy (jamesj):

48 = 8 x 6 = 2x2x2 x 2x3 Now the prime factorization of 32 is 32 = 2x2x2x2x2

OpenStudy (anonymous):

i dont know 2 and 3

OpenStudy (jamesj):

So comparing factorizations as we have done for the last two problems, 48 = 2 x 2 x 2 x 2 x x 3 32 = 2 x 2 x 2 x 2 x 2 ------------------------ gcf = So what are the prime factors in common across 48 and 32?

OpenStudy (jamesj):

Here's the last example again, 24 = 2 x 2 x 2 x 3 72 = 2 x 2 x 2 x 3 x 3 ------------------- GCF = 2 x 2 x 2 x 3 = 8 x 3 = 24

OpenStudy (jamesj):

So, what's the gcf of 32 and 48? 48 = 2 x 2 x 2 x 2 x x 3 32 = 2 x 2 x 2 x 2 x 2 ------------------------ gcf = What are the prime factors in common across 48 and 32?

OpenStudy (anonymous):

\[48=2x2x2x2 and \] im confused

OpenStudy (jamesj):

48 = 2 x 2 x 2 x 2 x x 3 32 = 2 x 2 x 2 x 2 x 2 ------------------------ gcf = 2 x 2 x 2 x 2 , yes and 2 x 2 x 2 x 2 = 4 x 4 = 16. hence the gcf of 48 and 32 is 16.

OpenStudy (jamesj):

So look carefully at the three examples. GCF of 45 and 60: 45 = 3 * 3 * 5 60 = 2 * 2 * 3 * 5 ------------------ GCF= 3 * 5 = 15 ============ GCF of 24 and 72: 24 = 2 x 2 x 2 x 3 72 = 2 x 2 x 2 x 3 x 3 ------------------- GCF = 2 x 2 x 2 x 3 = 8 x 3 = 24 ============ GCF of 48 and 32: 48 = 2 x 2 x 2 x 2 x x 3 32 = 2 x 2 x 2 x 2 x 2 ------------------------ gcf = 2 x 2 x 2 x 2 = 16

OpenStudy (jamesj):

See what we did in each example?

OpenStudy (anonymous):

so its 16 not 2 or 3 or yes

OpenStudy (anonymous):

what mean is yes

OpenStudy (jamesj):

Here are your remaining problems: GCF of (100,30) (54,36) (3,97,5) (4,20,24) (36,9,45) You do the next one by yourself,really. Step 1, find the prime factorization of 100 and 30. Go.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im doing this the fast way 10x10=100